OpenStudy (anonymous):

Consider the function f(x) = (x^2-9)/(x-4) . Determine all asymptotes of this function including horizontal, vertical, and oblique (slant).

OpenStudy (anonymous):

the degree of the numerator is 2 and the degree of the denominator is 1, so there is no vertical asymptote, but there is a slant asymptote because the degree of the numerator is larger by 1

OpenStudy (anonymous):

the vertical asymptote is easiest, it is where the function is undefined, where the denominator is 0 you can solve that with your eyeball right?

OpenStudy (anonymous):

you find the "slant asymptote" by division, i.e. divide \(x^2-9\) by \(x-4\) you can use synthetic division if you know it, otherwise you have to use polynomial long division, which is harder

OpenStudy (anonymous):

how do you feel about synthetic division?

OpenStudy (anonymous):

I think i'm fine with synthetic division, thanks a lot!

OpenStudy (anonymous):

when you do it you can ignore the remainder, because it is only the line that you need

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Sorry, one more question when I did synthetic division I got the numbers 1 4 25. Is that right..?

OpenStudy (anonymous):

And how would you find horizontal asymptote?

OpenStudy (anonymous):

ok there is NO horizontal asymptote because the degree of the numerator is larger than the degree of the denominator that is why you are finding the slant asymptote

OpenStudy (anonymous):

1 4 25 is wrong actually, it should be 1 4 7 but in fact is doesn't matter because this means \[\frac{x^2-9}{x-4}=x+4+\frac{7}{x-4}\] but the remainder is unimportant

OpenStudy (anonymous):

And there is also no vertical asymptote just by looking.

OpenStudy (anonymous):

the line \(y=x+4\) is your slant asymptote

OpenStudy (anonymous):

forget the remainder part, that is just the remainder

OpenStudy (anonymous):

oh yes indeed there is a vertical asymptote you get it by setting the denominator equal to zero and solve

OpenStudy (anonymous):

\[x-4=0\iff x=4\] and so your vertical asymptote is the vertical line \(x=4\)

OpenStudy (anonymous):

you good with this?

OpenStudy (anonymous):

Yeah I actually think I am, thanks again.